# Reservoir-Computing Model for Mapping and Forecasting Neuronal Interactions from Electrophysiological Data

### 23 Jan 2024 v2

@@ Acknowledgements
This work has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101033260 (project ISLAND) and the European Research Council (ERC) grant agreement No 788793 (project BACKUP). We would also like to thank Prof. Paolo Bettotti (University of Trento) for his support in simulations and computing infrastructures, Dr. Beatrice Vignoli (University of Trento) for her expert advisory in the biological settings, Prof. Michela Chiappalone (University of Genoa) for the training of MEA measurements and data analysis.
@@ I Experimental procedures
In some experiments, we employed optical stimulation to activate ChR2-expressing neurons. This method enables a more precise and localized manipulation of neural activity, stimulating specific regions within the network, as opposed to the broader influence of electrical stimulation. As the light source we used a Digital Light Processor (DLP) system. The system is a DLP E4500, which includes 3 LEDs, optics, a WXGA DMD (Wide Extended Graphics Array Digital Micro-mirror Device) and a driver board. The light engine can produce approximately 150 lumen at 15W LED power consumption. The blue LED (488nm) which is used in this work has a power of 600mW. The light from the LEDs impinges on the DMD which has 1039680 mirrors arranged in 912 columns by 1140 rows with a diamond array configuration. Each of these mirrors has two main possible inclinations that reflect the light in a different direction. This system allows to get patterned illumination with pre-loaded and custom patterns that can be chosen through the DLP E4500 software. Moreover, these patters can be pulsed in time, with both an internal or external trigger, with a nominal precision down to $\mu$s. The system supports 1-, 2-, 3-, 4-, 5-, 6-, 7-, and 8-bit images with a 912 columns $\times$ 1140 rows resolution. These images are pixel accurate, meaning that each pixel corresponds to a micro-mirror on the DMD. The light coming from the DLP system is collimated and aligned to the optical path of the microscope from the rear port of the system. As can be seen in FIG. The light from the DLP is collected by a macro TAMRON 90mm AF2.5 objective and the light pattern is imaged on the sample plane, where the MEA chip is located, a by 10$\times$ objective, while passing through a dichroic mirror (Chroma T505lpxr-UF1) which acts like a high-pass filter, reflecting all the wavelengths smaller than 505nm.
The ChR2-infected culture could be in parallel imaged using a microscopy system, where the signal of the green fluorescent protein (GFP) expressed by the ChR2 infected cells, is transmitted through a dichroic filter and detected by a CMOS camera. The culture image allows to direct the desired light pattern from the DLP directly to the region of interest in the culture.
The light stimulation comprised 300 pulses of 20 milliseconds each, delivered as blue light at a frequency of 0.5 Hz (with 2-second intervals between pulses.)
@@ II Data Pre-Processing
Following MEA recordings, the raw data was exported from MCS software and analyzed using a custom code written in Python. First the raw signals recorded on MEA were digitally filtered with a band-pass Butterworth filter with cutoff frequencies 0.3 and 3 KHz. Spikes were detected using the PTSD algorithm [33], setting the differential threshold (DF) to 8, refractory period (RP) to 1ms a
After detecting spikes, we proceeded to apply a burst detection algorithm inspired by the approach outlined in Ref. [31], but with some minor modification. By utilizing the spike trains, we derived the inter-spike intervals (ISI) for each channel within the recording. Subsequently, data from all channels was aggregated into a single ISI histogram, constructed with fixed bins of $\Delta\log_{10}\{ISI\}=0.1$ (in units of $\log_{10}$(s). The ISI threshold was then calculated by the following algorithm:
@@
  • Sort maxima by their significance, where significance is determined by $S_{peak}=p\cdot w$; $p$ is the peak prominence and $w$ is the peak full-width at half-prominence (FWHP).
  • Choose the most significant peak within the bins between -3 and -2, corresponding to 1-10 $ms$, characterizing typical inter-burst intervals [31].
  • Select the most significant peak within the bins spanning from -3 to -2, which corresponds to the range of 1 to 10 milliseconds. This peak characterizes the typical fast intra-burst intervals (for each channel).
  • Choose the most significant peak within the bins greater than -2, corresponding to intervals longer than 10 milliseconds. This peak characterizes the typical inter-burst intervals (for each channel).
  • calculate the ISI threshold by:
@@
    where $x_{l},x_{r}$ are the bin value of the left and right peaks, respectively; $w_{l},w_{r}$ are the FWHP of the left and right peaks, respectively.
@@
In the subsequent phase, we examined instances where bursting activity spread across the culture, indicating periods when bursts were detected on multiple channels within a specific time frame - also known as Network Bursts (NBs) [43]. NB was identified if burst occurred on at least two distinct channels $j$ and $k$ with respective starting times, $t_{B}^{j}$ and $t_{B}^{k}$, were separated by a time interval no greater than $\langle L_{B}\rangle/2$, where $\langle L_{B}\rangle$ is the average length of a burst. Consequently, NB starting times $t_{NB}^{j}$ were registered.
@@ III Model Parametrization
For instance, in the lasso regression (Eq. 2.1), the optimal $\mathcal{W}_{out}$ was determined for each set of $\mathcal{W}{res}$, $\mathcal{W}_{in}$, and $\hat{\mathcal{S}}$. We specifically evaluated the stability of connections within the intrinsic connectivity matrix (ICM), $\mathcal{T}_{0}$ (Eq. 6). This matrix results from the product of $\mathcal{W}{out}$ and $\hat{\mathcal{S}}$ with $\mathcal{W}{in}$. We assessed the consistency of connections in the ICM across different training sessions, quantifying it as the ICM confidence as follows:
@@
Here, $\Gamma_{CM}$ represents the confidence value, $\sigma_{\mathcal{T}_{0}}$ stands for the standard deviation of the connection weights across the repetitions, and $\langle\mathcal{T}_{0}\rangle$ denotes the mean connection weights averaged over the repetitions. The assessment of response prediction stability, as outlined in Section 4.5.2, involved conducting the test with multiple initializations and calculating the mean response. The degree of stability is visualized by the error bars, which represent the standard deviation of the predicted response for each channel (Figs. 4(c) and 5(c) ,Section 4.5.2)
Regarding the hyperparameters $m$ and $\alpha$, we characterized their effect on model outcomes in all steps: Validation, Connectivity prediction and Response prediction. Fig. LABEL:fig:supp illustrates the characterization of some of the model’s metrics as a function of hyperparameters $m$ and $\alpha$.
@@ IV External Test Parameters (Stimulation Intensity)
During the test phase where we examined how accurate the model can predict the network response to a given input (stimulus), we found that it is useful to tune the stimulation intensity as a fitting parameter in order to minimize the error between the ground truth response $\mathbf{U}$ and the prediction $\tilde{\mathbf{U}}$ in terms of $\bar{R}$ metric (Eq. (9)). Assuming that the stimulation is applied around node (electrode) $j$, we represent the input state (the stimulation) as:
$\mathbf{u}_{in}=I\cdot\delta_{i,j}\;\;\;i=1,2,...,N_{ch}$ (33)
where $I$ is the intensity and $\delta_{i,j}$ is the Kronecker delta.
@@ V Performance Optimization
Figure 10: Parametrization of RC model estimators and metrics (dependence on hyperparameters $m$ and $\alpha$). All figures refer to results obtained from the same NEST simulation dataset presented in Figs 2 and 5. (a) validation loss (Eq. (4.3.2)). (b) Pearson correlation of the ICM with the ground truth connectivity matrix. (c) ROC AUC for the connectivity prediction. (d) Confidence measure of the ICM (Eq. (32)). (e) $\bar{R}$ metric of the response prediction. (f) ROC AUC metric for the response prediction.
Figure 11: Response test metrics dependence on input intensity of the stimulus. The intensity was used as a tuning parameter to fit the model response to the experimental. (a) $\bar{R}$ (b) The overall time lag between the modeled response and the experimental, $\bar{\tau}_{l}$ (Eq. (30)). (c) ROC AUC.
VI More Supplementary Figures
Figure 12: The principle of the XRMSE metric. (a) Ground truth and prediction data. (b) The outcome of XRMSE for the prediction of (a)- Eq. (26). The red point indicates the optimum point from which the channel error and the time lag ($\varepsilon_{i}$, $\tau_{l,i}$ Eq. (28)) are obtained. This procedure is done over all the channels of the network, followed by the total error, $\bar{R}$, and time lag, $\bar{\tau}_{l}$, estimation (Eqs. (29), (30)).

### 05 Mar 2024 v3

@@ Acknowledgements
This work has received funding from European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No 101033260 (project ISLAND) and the European Research Council (ERC) grant agreement No 788793 (project BACKUP). We also acknowledge Fondazione Caritro for their additional funding.
We would also like to thank Prof. Paolo Bettotti (University of Trento) for his support in simulations and computing infrastructures, Dr. Beatrice Vignoli (University of Trento) for her expert advice in the biological settings, Prof. Michela Chiappalone (University of Genoa) for the training in MEA measurements and data analysis.
Supplementary Figures
Figure Supplementary 1: Enhanced visual representation illustrating the interconnection between the computational circuitry of the RC model (Fig. 0(c)) and the intricate circuitry of biological signals.
Figure Supplementary 2: Illustration of the stimulus test designed to assess the RC model’s network response prediction performance to a local stimulus. On the left, an electrode map highlights the stimulated electrode (in red) and the responsive electrodes (in yellow), demonstrating spatial prediction. This is further evaluated against the experimental/simulated response using the ROC AUC metric. On the right, a detailed temporal response from all involved electrodes showcases spatio-temporal prediction. This is quantitatively compared to the experimental/simulation test using the $\bar{R}$ metric (refer to Eq. (9)).
Figure Supplementary 3: Response test (NEST simulation data). A detailed map of each channel’s response to a stimulus is given at channel 11 (designated in yellow star). The time scale (X-axis) is given in the legend. Each point corresponds to a time step of 2ms. The $y$ axis is the normalized ISR in arbitrary units, scaled identically. $\bar{R}=0.158$, $\bar{\tau}_{l}=0.55ms$ at $\alpha=0.5$ and $m=50$.
((a))
((b))
((c))
((d))
((e))
((f))
Figure Supplementary 4: Parametrization of RC model estimators and metrics (dependence on hyperparameters $m$ and $\alpha$). All figures refer to results obtained from the same NEST simulation dataset presented in Figs 2 and 5. (a) validation loss (Eq. (5.3.2)). (b) Pearson correlation of the ICM with the ground-truth connectivity matrix. (c) ROC AUC for the connectivity prediction. (d) Confidence measure of the ICM (Eq. (32)). (e) $\bar{R}$ metric of the response prediction. (f) ROC AUC metric for the response prediction.
((a))
((b))
((c))
((d))
Figure Supplementary 5: Different network dynamics derived from electrophysiological data (taken from NEST simulation). (a) Raster plot illustrating the simulated electrophysiological activity of a 60-population in-silico network. (b) An example of Inter-Spike-Interval (ISI) histogram from a single population, recorded on one channel, revealing a bimodal shape indicative of bursting activity, with a prominent peak at shorter time scales ($10^{0}-10^{2}$ ms). (c) Aggregated ISI histogram representing the entire network’s ISIs across all channels (of intervals taken on a single channel), exhibiting a typical bimodal shape with a major peak in the $10^{-1}-10^{1}$ ms range and a secondary peak in the $10^{2}-10^{3}$ ms range, signifying modulated bursting activity. Intra-burst intervals are characterized by the left peak, while inter-burst intervals are marked by the right peak. The red star indicates the ISI threshold ($ISI_{th}$) determined using the burst detection algorithm in Sec. II.a. (d) Inter-Burst-Intervals (IBI) histogram showcasing the distribution of IBIs between different channels, depicting typical network bursting (NB) behavior. This histogram aids in deriving training data, divided into batches of NBs (Sec. 5.3). The red star denotes the estimated integration time (in this case 2 ms), by the algorithm described in Sec. II.b.
((a))
((b))
((c))
Figure Supplementary 6: Response test metrics dependence on input intensity of the stimulus. The intensity was used as a tuning parameter to fit the model response to the experiment. (a) $\bar{R}$ (b) The overall time lag between the modeled response and the experiment, $\bar{\tau}_{l}$ (Eq. (30)). (c) ROC AUC.
((a))
((b))
Figure Supplementary 7: The principle of the cross-root-mean- square-error ($R$) metric. (a) ground-truth and prediction data. (b) The outcome of $R$ for the prediction of (a)- Eq. (26). The minima describes the channel error and the time lag ($\varepsilon_{i}$, $\tau_{l,i}$ Eq. (28)) are obtained. This procedure is done over all the nodes of the network, followed by the total error, $\bar{R}$, and time lag, $\bar{\tau}_{l}$, estimation (Eqs. (29), (30)).
@@ I Experimental procedures
In some experiments, we employed optical stimulation to activate ChR2-expressing neurons. This method enables more precise and localized manipulation of neural activity, stimulating specific regions within the network, as opposed to the broader influence of electrical stimulation. As the light source, we used a Digital Light Processor (DLP) system. The system is a DLP E4500, which includes 3 LEDs, optics, a WXGA DMD (Wide Extended Graphics Array Digital Micro-mirror Device), and a driver board. The blue LED (488nm) which is used in this work has a power of 600mW. The light from the LEDs impinges on the DMD which has 1039680 mirrors arranged in 912 columns by 1140 rows with a diamond array configuration. This system allows patterned illumination with pre-loaded and custom patterns that can be chosen through the DLP E4500 software. Moreover, these patterns can be pulsed in time, with both an internal or external trigger, with a nominal precision down to $\mu$s. The system supports 1-, 2-, 3-, 4-, 5-, 6-, 7-, and 8-bit pixel resolution images with a 912 columns $\times$ 1140 rows resolution, meaning that each pixel corresponds to a micro-mirror on the DMD. The light coming from the DLP system is collimated and aligned to the optical path of the microscope from the rear port of the system. It is then collected by a macro TAMRON 90mm AF2.5 objective and the light pattern is imaged on the sample plane, where the MEA chip is located, a by 10$\times$ objective, while passing through a dichroic mirror (Chroma T505lpxr-UF1) which acts like a high-pass filter, reflecting all the wavelengths smaller than 505nm.
The ChR2-infected culture could be imaged using a microscopy system, where the signal of the green fluorescent protein (E-YFP) expressed by the ChR2-infected cells, is transmitted through a dichroic filter and detected by a CMOS camera. The culture image allows to direct the desired light pattern from the DLP directly to the region of interest in the culture.
[…]
@@ II Data Pre-Processing
Following MEA recordings, the raw data was exported from MCS software and analyzed using a custom code written in Python. First, the raw signals recorded on MEA were digitally filtered with a band-pass Butterworth filter with cutoff frequencies of 0.3 and 3 KHz. Spikes were detected using the PTSD algorithm [38], setting the differential threshold (DF) to 8, the refractory period (RP) to 1 ms, and the peak lifetime period (PLP) to 2 ms.
After detecting spikes, we proceeded to apply a burst detection algorithm inspired by the approach outlined in [36], but with some minor modifications. By utilizing the spike trains, we derived the inter-spike intervals (ISI) for each channel within the recording. Subsequently, data from all channels was aggregated into a single ISI histogram (Fig. 4(c)), constructed with fixed bins of $\Delta\log_{10}\{ISI\}=0.1$ (in units of $\log_{10}$(ms). The ISI threshold was then calculated by the following algorithm:
@@
  • Sort maxima by their significance, where significance is determined by $S_{peak}=p\cdot w$; $p$ is the peak prominence and $w$ is the peak full-width at half-prominence (FWHP), expressed in $\log_{10}(ms)$ units.
  • Choose the most significant peak within the bins between $1$ and $10$ ms, characterizing typical inter-burst intervals [36].
  • Select the most significant peak within the bins spanning from $1$ to $10$ ms. This peak characterizes the typical fast intra-burst intervals (for each channel).
  • Choose the most significant peak within the bins greater than $10$ ms. This peak characterizes the typical inter-burst intervals (for each channel).
  • calculate the ISI threshold (in ms) by:
@@
    where $x_{l},x_{r}$ are the bin value of the left and right peaks, respectively; $w_{l},w_{r}$ are the FWHP of the left and right peaks, respectively- all expressed in $\log_{10}(ms)$ units.
@@
In the subsequent phase, we examined instances where bursting activity spread across the culture, indicating periods when bursts were detected on multiple channels within a specific time frame - also known as Network Bursts (NBs) [36, 47, 48]. NB was identified if bursts occurred on at least two distinct channels $j$ and $k$ with respective starting times, $t_{B}^{j}$ and $t_{B}^{k}$, were separated by a time interval no greater than $\langle L_{B}\rangle/2$, where $\langle L_{B}\rangle$ is the average length of a burst. Consequently, NB starting times $t_{NB}^{j}$ were registered, followed by the procedure described in Sec. 5.3.2.
II.b Estimation of the Integration Time
We define integration time as the duration required for each channel (neuronal circuit) to assimilate incoming information and transmit a response through the network. A detailed analysis of bursting dynamics, represented by the Inter-Burst-Interval (IBI) histogram showcasing the distribution of IBIs between distinct channels (refer to Fig. 4(d)), reveals a prominent peak. This peak represents a characteristic range of time intervals during which one channel responds to the burst of another. Hence, we deduce that information flows within this specific rate, leading us to designate it as the effective integration time.
Consequently, we utilize this characteristic time to sample the state $\mathbf{y}$ of the network at time steps corresponding to the integration time. Subsequently, we bin the spike trains into intervals aligning with this time, generating the Instantaneous Spike Rate (ISR) traces (see Sec. 5.3).
Accordingly, we implemented the following algorithm for estimating the integration time:
  • Detect local maxima in the IBI histogram.
  • Sort maxima by their significance, where significance is determined by $S_{peak}=p\cdot w$; $p$ is the peak prominence and $w$ is the peak full-width at half-prominence (FWHP), expressed in $\log_{10}(ms)$ units.
  • If the most significant peak is found at a value less than 2 ms $\rightarrow$ set 2 ms as the integration time. This assumption arises from the rapid dynamics, and typically, this interval can accommodate at most one or two spikes.
  • If no peaks are found in this range $\rightarrow$ set 5 ms as the integration time. This pertains specifically to slow dynamics; unless the network’s dynamics are rich within this regime, poor performance is expected.
  • Otherwise $\rightarrow$ the most significant peak in this range indicates the integration time, with its value is rounded to the closest 0.1ms multiple (for usage of rounded numbers in the binning step).
Through empirical testing of the RC model’s performance across various integration times, we demonstrated that the algorithm consistently delivers close-to-optimal results in a substantial number of cases.
@@ III Model Parametrization
For instance, in the lasso regression (Eq. 2), the optimal $\mathcal{W}_{out}$ was determined for each set of $\mathcal{W}_{res}$, $\mathcal{W}_{in}$, and $\hat{\mathcal{S}}$. We specifically evaluated the stability of connections within the intrinsic connectivity matrix (ICM), $\mathcal{T}_{0}$ (Eq. 7). This matrix results from the product of $\mathcal{W}_{out}$ and $\hat{\mathcal{S}}$ with $\mathcal{W}_{in}$. We assessed the consistency of connections in the ICM across different training sessions, quantifying it as the ICM confidence as follows:
@@
Here, $\Gamma_{CM}$ represents the confidence value, $\sigma_{\mathcal{T}_{0}}$ stands for the standard deviation of the connection weights across the repetitions, and $\langle\mathcal{T}_{0}\rangle$ denotes the mean connection weights averaged over the repetitions. The assessment of response prediction stability, as outlined in Section 5.5.2, involved conducting the test with multiple initializations and calculating the mean response. The degree of stability is visualized by the error bars, which represent the standard deviation of the predicted response for each channel (Figs. Supplementary 3 and 6(c)).
Regarding the hyperparameters $m$ and $\alpha$, we characterized their effect on model outcomes in all steps: Validation, Connectivity prediction, and Response prediction. Fig. Supplementary 4 illustrates the characterization of some of the model’s metrics as a function of hyperparameters $m$ and $\alpha$.
@@ IV External Test Parameters (Stimulation Intensity)
During the test phase where we examined how accurately the model can predict the network response to a given input (stimulus), we found that it is useful to tune the stimulation intensity as a fitting parameter to minimize the error between the ground-truth response $\mathbf{U}$ and the prediction $\tilde{\mathbf{U}}$ in terms of $\bar{R}$ metric (Eq. (9)). Assuming that the stimulation is applied around node (electrode) $p$, we represent the input state (the stimulation) as:
$\mathbf{u}_{in}=I\cdot\delta_{i,p}\;\;\;i=1,2,...,N_{ch}$ (33)
where $I$ is the intensity and $\delta_{i,p}=0$, for $i\neq p$ and $\delta_{i,p}=1$ for $i=p$ (Kronecker delta).